find all the 5th roots of 5-i
1
Expert's answer
2019-12-17T05:08:43-0500

What is actually asked here is to solve the equation


"z^5=5-i."

In other words, the modulus here is


"m=\sqrt{5^2+(-1)^2}=\sqrt{26},"

and the argument is


"\phi=\text{atan}\frac{-1}{5}=-0.197\text{ rad}."

Therefore, we can rewrite the equation into the following polar form:


"z^5=\sqrt{26}\cdot e^{-0.197i+2k\pi i}."

Simplify:

"z^5=\sqrt{26}\cdot e^{(-0.197+2k\pi)i}."

Let


"z=re^{i\theta}."

Hence:


"\big(re^{i\theta}\big)^5=\sqrt{26}\cdot e^{(-0.197+2k\pi)i},\\
\space\\
r^5e^{5i\theta}=\sqrt{26}\cdot e^{(-0.197+2k\pi)i},"

equate the modulus and the power of "e":


"r^5=\sqrt{26},\space\space\space\space\space\space\space5\theta=-0.197+2k\pi;\\
\space\\
r=\sqrt[10]{26},\space\space\space\space\space\space\space\theta=\frac{2k\pi-0.197}{5}.\\"

Now the question is: what values of "k" must we choose in order to get five different and unique roots? The answer is the values starting from zero to "n-1", that is,


"k=0,1,2,3,4."

Substitute these values and get 5 roots:


"z_1=\sqrt[10]{26}\cdot e^\frac{-0.197}{5},\\
z_2=\sqrt[10]{26}\cdot e^\frac{2\pi-0.197}{5},\\
z_3=\sqrt[10]{26}\cdot e^\frac{4\pi-0.197}{5},\\
z_4=\sqrt[10]{26}\cdot e^\frac{6\pi-0.197}{5},\\
z_5=\sqrt[10]{26}\cdot e^\frac{8\pi-0.197}{5}."